# Improving Generalization of Universal Adversarial Perturbation via Dynamic Maximin Optimization

### v1

@@ Abstract
Deep neural networks (DNNs) are susceptible to universal adversarial perturbations (UAPs). These perturbations are meticulously designed to fool the target model universally across all sample classes. Unlike instance-specific adversarial examples (AEs), generating UAPs is more complex because they must be generalized across a wide range of data samples and models. Our research reveals that existing universal attack methods, which optimize UAPs using DNNs with static model parameter snapshots, do not fully leverage the potential of DNNs to generate more effective UAPs. Rather than optimizing UAPs against static DNN models with a fixed training set, we suggest using dynamic model-data pairs to generate UAPs. In particular, we introduce a dynamic maximin optimization strategy, aiming to optimize the UAP across a variety of optimal model-data pairs. We term this approach DM-UAP. DM-UAP utilizes an iterative max-min-min optimization framework that refines the model-data pairs, coupled with a curriculum UAP learning algorithm to examine the combined space of model parameters and data thoroughly. Comprehensive experiments on the ImageNet dataset demonstrate that the proposed DM-UAP markedly enhances both cross-sample universality and cross-model transferability of UAPs. Using only 500 samples for UAP generation, DM-UAP outperforms the state-of-the-art approach with an average increase in fooling ratio of 12.108%.
Codehttps://github.com/yechao-zhang/DM-UAP
@@ 1 Introduction
Adding to this concern is the emergence of a notable class of adversarial techniques known as universal adversarial perturbation (UAP), which are designed to induce mispredictions against universal input samples (Moosavi-Dezfooli et al. 2017; Khrulkov and Oseledets 2018; Poursaeed et al. 2018; Shafahi et al. 2020). In addition to this characteristic of cross-sample universality, UAPs could also possess another element of generalization in adversarial attacks, termed cross-model transferability. This denotes the capacity of (AEs) to preserve their efficacy across different DNN models (Hu et al. 2022; Li et al. 2024). Together, these capabilities enable UAPs to facilitate the generation of transferable AEs at scale, compromising various DNN architectures and multiple machine learning tasks simultaneously (Zhong and Deng 2022; Xie et al. 2021; Ding et al. 2021; Wang et al. 2023; Zhou et al. 2023a, b, 2024). Consequently, UAPs have garnered significant attention and have become a focal point of extensive research in the field.
@@ 4 Experiment
Method ANet GNet V16 V19 R152 Avg
@@ 4.1 Generalization Performance of UAPs
Method Samples ANet V16 GNet V19 R50 R152 Avg
@@ 4.2 Scalability Performance of UAPs
Table 3: The fooling ratio (%) on five models in the black-box setting by different UAP attack methods. The UAPs are crafted on AlexNet, GoogleNet, VGG16, VGG19, and ResNet152, respectively. We conduct the evaluation using 10,000 and 500 images for training, respectively. The respective last column represents the average improvement or deterioration.
@@
Figure 2: Average fooling ratio (%) on five models in the diverse-sample training scenarios, using VGG19.
@@
Model UAP Method ANet GNet V16 V19 R152 Avg
@@
Figure 3: Ablation study on model and data optimization. “M” means “Model Optimization” and “D” means “Data Optimization”, “M-D” denotes “Model-Data Optimization”, “D-M” denote “Data-Model Optimization”. (a) Fooling ratios of DM-UAP with/without curriculum learning, i.e, increasing neighborhood sizes. (b) Fooling ratio by different optimization orders in white-box setting for five models. (c) Fooling ratios of DM-UAP with/without data optimization for different model neighborhood sizes. (d) Fooling ratios of DM-UAP with/without model optimization for different data neighborhood sizes.
@@ 5 Limitation and Future Work
Method Time consumption Memory consumption Average attack success rate(%)

### v3

@@ Abstract
Deep neural networks (DNNs) are susceptible to universal adversarial perturbations (UAPs). These perturbations are meticulously designed to fool the target model universally across all sample classes. Unlike instance-specific adversarial examples (AEs), generating UAPs is more complex because they must be generalized across a wide range of data samples and models. Our research reveals that existing universal attack methods, which optimize UAPs using DNNs with static model parameter snapshots, do not fully leverage the potential of DNNs to generate more effective UAPs. Rather than optimizing UAPs against static DNN models with a fixed training set, we suggest using dynamic model-data pairs to generate UAPs. In particular, we introduce a dynamic maximin optimization strategy, aiming to optimize the UAP across a variety of optimal model-data pairs. We term this approach DM-UAP. DM-UAP utilizes an iterative max-min-min optimization framework that refines the model-data pairs, coupled with a curriculum UAP learning algorithm to examine the combined space of model parameters and data thoroughly. Comprehensive experiments on the ImageNet dataset demonstrate that the proposed DM-UAP markedly enhances both cross-sample universality and cross-model transferability of UAPs. Using only 500 samples for UAP generation, DM-UAP outperforms the state-of-the-art approach with an average increase in fooling ratio of 12.108%. Our codes are publicly available at https://github.com/yechao-zhang/DM-UAP
@@ 1 Introduction
Adding to this concern is the emergence of a notable class of adversarial techniques known as universal adversarial perturbation (UAP), which are designed to induce mispredictions against universal input samples (Moosavi-Dezfooli et al. 2017; Khrulkov and Oseledets 2018; Poursaeed et al. 2018; Shafahi et al. 2020). In addition to this characteristic of cross-sample universality, UAPs could also possess another element of generalization in adversarial attacks, termed cross-model transferability. This denotes the capacity of (AEs) to preserve their efficacy across different DNN models (Hu et al. 2022; Zhang et al. 2024; Zhou et al. 2023). Together, these capabilities enable UAPs to facilitate the generation of transferable AEs at scale, compromising various DNN architectures and multiple machine learning tasks simultaneously (Zhong and Deng 2022; Xie et al. 2021; Ding et al. 2021; Wang et al. 2023). Consequently, UAPs have garnered significant attention and have become a focal point of extensive research in the field.
@@ 4 Experiment
Method AlexNet GoogleNet VGG16 VGG19 ResNet152 Average
@@ 4.1 Generalization Performance of UAPs
Method Samples AlexNet VGG16 GoogleNet VGG19 ResNet50 ResNet152 Average
@@ 4.2 Scalability Performance of UAPs
Table 3: The fooling ratio (%) on five models in the black-box setting by different UAP attack methods. The UAPs are crafted on AlexNet, GoogleNet, VGG16, VGG19, and ResNet152, respectively. We conduct the evaluation using 10,000 and 500 images for training, respectively. The average improvement or deterioration of AT-UAP, SGA, and DM-UAP are also provided, with improvements highlighted in green and deteriorations in red. * indicates the white-box model.
@@
Figure 2: Average fooling ratio (%) on five models in the diverse-sample training scenarios. The UAPs are crafted by UAP, GAP, SPGD, and SGA, and our DM-UAP on VGG19.
@@
Model UAP Method AlexNet VGG16 GoogleNet VGG19 ResNet152 Average
@@
Table 5: The fooling ratio (%) in the attack-under-defense setting on five models by different UAP generation methods.
Defense UAP Method AlexNet VGG16 GoogleNet VGG19 ResNet152 Average
JPEG SPGD 91.9 79.6 64.6 60.0 42.6 67.7
SGA 91.3 82.2 81.5 81.3 56.6 78.6${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+10.9)}}}$
Ours 93.0 79.5 89.6 84.6 53.7 80.1${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+12.36)}}}$
NRP SPGD 61.2 40.7 56.3 52.9 35.8 49.4
SGA 61.8 40.1 55.1 55.1 36.5 49.7${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+0.3)}}}$
Ours 63.1 41.3 56.1 55.3 38.6 50.9${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+1.5)}}}$
Diffpure SPGD 36.0 38.2 36.7 38.2 39.7 37.8
SGA 39.4 37.5 37.9 36.8 37.4 37.8${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+0)}}}$
Ours 38.5 38.8 39.3 40.8 41.0 39.7${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+1.9)}}}$
@@
Figure 3: Ablation study on model and data optimization. (a) Fooling ratios of DM-UAP with/without curriculum learning, i.e, increasing neighborhood sizes. (b) Fooling ratio by different optimization orders in white-box setting for five models. (c) Fooling ratios of DM-UAP with/without data optimization for different model neighborhood sizes. (d) Fooling ratios of DM-UAP with/without model optimization for different data neighborhood sizes.
Attack-under-Defense Setting. We evaluate the attack performance of AEs crafted by different UAP methods under three common defenses: JPEG (Das et al. 2018), NPR (Naseer et al. 2020), and DiffPure(Nie et al. 2022). JPEG uses lossy compression to remove adversarial perturbations. NPR trains a purifier network that minimizes the difference in perceptual features between clean and adversarial images. DiffPure uses Gaussian noise to smooth out adversarial perturbations and then denoises images using a pre-trained diffusion model. Tab. 5 shows that DM-UAP performs the best on average under every defense.
@@ 5 Limitation and Future Work
UAP method Time consumption Memory consumption Average attack success rate(%)